A 2.0-kg particle has an initial velocity of (Si-4) m/s, Sometime later, its velocity is 7i +3j assuming no energy is lost in the process? m/s. How much work was done by the resultant force during this time interval, b. 49 J c. 19 d. 53 J e 27
Answer #1

work done = change in KE

W = \Delta KE

W = 1/2 mvf2 - 1/2 mvi2

vf = √(72+32) = √58

vi = √(52+42) = √41

W = (1/2 )( 2kg) (58-41) = 17 J (a)
Answer #2

we have m= 2kg  and  v= 5i - 4j  and  vf= 7i + 3j

W = △KE = 1/2 mv2

W = 1/2 mvf2 - 1/2 mvi2

vf = √((7)2+(3)2) = √58

vi = √((5)2+(-4)2) = √41

W = (1/2 )( 2 )(58)2 - (1/2)(2)(41)2 17 J

answer = 17 J (a)


answered by: Sarah Hazem